Tutors
Simplify
log3 64/log3 16 the threes are smaller and placed below the log
thank you for your help

Answer by nabla(475)

By the properties of logarithms, we can rewrite that as:
log16 64
Which is the same as:
16^x=64
2^(4x)=2^6
So, 4x=6
x=3/2